OpenStudy (anonymous):

T(t)=20+5cos(pi/12(t-3)) where T is the temperature . When was temp below 17.5C When graphing time is on the x axis and temp on y?

OpenStudy (perl):

you can graph T(t)=20+5cos(pi/12(t-3)) and T = 17.5 and see where they intersect

OpenStudy (anonymous):

this is a weird question or I am missing something but temp at noon is 23.5 and by calculations it say min will occur at 15 and max at 3 bizarre, shouldn't it be the other way around?

OpenStudy (perl):

yes that is odd

OpenStudy (perl):

wait, is T temperature since noon?

OpenStudy (anonymous):

i have checked and checked

OpenStudy (perl):

t= 0 corresponds to noon temperature

OpenStudy (perl):

can you also post the directions to the last question , that would help

OpenStudy (anonymous):

ahhhhhhh yes, thank you, thought I was going nutty, should read the question properly LOL

OpenStudy (perl):

yes that makes sense, min occurs at t = 15 hours after noon, and max occurs t = 3 hours after noon

OpenStudy (perl):

now for the question above

OpenStudy (perl):

can i send you a link

OpenStudy (perl):

https://www.desmos.com/screenshot/lusqpaxs9t

OpenStudy (anonymous):

so it is best if I draw a line through at 17.5?

OpenStudy (anonymous):

That would be excellent or can I just do it on desmos?

OpenStudy (perl):

yes

OpenStudy (anonymous):

cool :)

OpenStudy (perl):

you have to zoom out though https://www.desmos.com/calculator/o3gakiua8i

OpenStudy (perl):

scroll down to zoom out

OpenStudy (anonymous):

\[f(x)=2-\sqrt{x/2+3}\] How do I work out the range of this function?

OpenStudy (perl):

the range of y= sqrt(x) is [0,oo) the range of y = -sqrt(x) is (-oo,0] then add to two that (-oo, 2]

OpenStudy (perl):

alternatively you can graph the function into desmos

OpenStudy (anonymous):

thanks again :)

OpenStudy (anonymous):

Are you still there?